OpenStudy (anonymous):

Write the rule for each sequence in words. Then find the next three terms. 35, 70, 105, 140. . .

OpenStudy (raden):

1st 35 (known) 2nd is 70 = 35 + 35 3rd is 105 = 70 + 35 4th is 140 = 105 + 35 ... ... and so on does that make sense ?